OpenStudy (anonymous):

solve cosx cot x over 1 - sinx = 3

OpenStudy (anonymous):

cosx * cosx)/sinx /(1-sin x)=3 cos^2 x/{sinx(1-sinx)}=3 (1-sin^2 x)/{sinx(1-sinx)}=3 (1+sinx)(1-sinx)//{sinx(1-sinx)}=3 cancel (1-sinx) (1+sinx)/sinx=3 divide by sinx 1/sinx=2 sinx=1/2 x=30 degrees